Suppose Mikes utility function for apples and bananas is U(A, B) = AB. What is the marginal utility of apples?* Your answer What is the marginal utility of bananas?* Your answer What is the marginal rate of substitution for apples with 2 bananas? * Your answer
Answer #1

1.

Marginal utility of apples = MUa

MUa = d(A^2B)/dA

MUa = B*2A + A^2*0

MUa = 2AB

B.

Marginal utility of bananas = MUb

MUb = d(A^2B)/db

MUb = A^2 + B*0

MUb = A^2

C.

Marginal rate of substitution= MRSab

MRSab = MUa/MUb = 2AB/A^2

MRSab = 2B/A